OpenStudy (anonymous):

plz help What are the real or imaginary solutions of the polynomial equation? 27x^3+125=0 I don't just want the answer please put an explanation with it.

OpenStudy (loser66):

27x^3 = (3x)^3 125 = 5^3 apply a^3 +b^3 = (a+b) (a^2 -ab+b^2)